OpenStudy (anonymous):

log5(4)+log5(2x)=log(5)24

OpenStudy (kropot72):

\[\log_{5} 4+\log_{5} 2x=\log_{5} 24\]

OpenStudy (callisto):

Is the question like this? \[log_5(4)+log_5(2x)=log_524\]

OpenStudy (callisto):

If it is, then \[log_5(4)+log_5(2x)=log_524\]\[log_5(4 \times2x)=log_524\]\[4(2x)=24\] I think you can solve it :)

OpenStudy (anonymous):

@Callisto does x=10

OpenStudy (callisto):

no.... 4(2x) = 24 8x = 24 x = ... ?

OpenStudy (anonymous):

oh. x=3

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

what if the = does not have log

OpenStudy (anonymous):

like this log10(4)+log10 (w)=2

OpenStudy (callisto):

hmm... \[2=log_{10}10 = log_{10}10^2\]

OpenStudy (anonymous):

no, the example i showed you

OpenStudy (callisto):

So, \[log_{10}(4)+log_{10} (w)=2\]\[log_{10}(4)+log_{10} (w)=log_{10}10^2\]\[log_{10}(4w)=log_{10}10^2\]\[4w = 10^2\]Solve w

OpenStudy (anonymous):

oh so if the 2 is a 4 it would be 10^4

OpenStudy (callisto):

If the base is still 10, then yes :)

OpenStudy (anonymous):

k, thx so, if the bas ewas 6 it would be 6^2

OpenStudy (callisto):

Yes!~

OpenStudy (anonymous):

k, thx bro.

OpenStudy (callisto):

Welcome :) it's a sis though lol

OpenStudy (anonymous):

oh sorry sis! :)